OpenStudy (anonymous):

If an atom has 18 protons, 22 neutrons, and 15 electrons, what is the atom's electrical charge? a.-7 b. +3 c. -4 d. +4

OpenStudy (anonymous):

There has to be even amount of protons, neutrons and electrons or else its called an "ion" Nuetrons are neutral. Protons are positively charged. Electrons are negatively charged. If number in those 3 are unbalanced (gained or lost) it would be an ion.

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20090324203606AAr3Fwt read this. this will answer your question more.

OpenStudy (anonymous):

here are your options a.-7 b. +3 c. -4 d. +4

OpenStudy (anonymous):

(Atomic number) - (Number of electrons present) = (overall electrical charge of atom) http://answers.yahoo.com/question/index?qid=20080829165423AAP4S9q

OpenStudy (anonymous):

atomic number is how many protons are present. \[p + e = ec\]

OpenStudy (anonymous):

18 - 15 = 3 The answer is +3.